OpenStudy (anonymous):

Please help I will given metal given the triangle a=27 b=46 a=71degrees find B

OpenStudy (johnweldon1993):

You can use the law of sines for this... \[\large \frac{\sin 71}{27} = \frac{\sin(B)}{46}\] And solve for B

OpenStudy (anonymous):

Did you get 87.2?

OpenStudy (johnweldon1993):

Hang on...is it SIDE A = 27 and SIDE B = 46, ANGLE A = 71 and we need ANGLE B ?

OpenStudy (anonymous):

Yes

OpenStudy (johnweldon1993):

Okay...then this triangle does not exist? \[\large \frac{\sin71}{27} \times 46 = 1.61\] \[\large \sin^{-1} 1.61 = \cancel{\text{Error}}\]